# Neural Incremental Input-to-State Stable Control Lyapunov Functions for Unknown Continuous-time Systems

### v1

@@ Abstract
This work primarily focuses on synthesizing a controller that guarantees an unknown continuous-time system to be incrementally input-to-state stable ($\delta$-ISS). In this context, the notion of $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for the continuous-time system is introduced. Combined with the controller, the $\delta$-ISS-CLF guarantees that the system is incrementally stable. As the paper deals with unknown dynamical systems, the controller as well as the $\delta$-ISS-CLF are parametrized using neural networks. The data set used to train the neural networks is generated from the state space of the system by proper sampling. Now, to give a formal guarantee that the controller makes the system incrementally stable, we develop a validity condition by having some Lipschitz continuity assumptions and incorporate the condition into the training framework to ensure a provable correctness guarantee at the end of the training process. Finally, we demonstrate the effectiveness of the proposed approach through several case studies: a scalar system with a non-affine, non-polynomial structure, a one-link manipulator system, a nonlinear Moore-Greitzer model of a jet engine, and a rotating rigid spacecraft model.
@@ 1 Introduction
Input-to-state stability has been a promising tool for analyzing and studying robust stability questions for nonlinear control systems. This type of analysis, which comes under the domain of traditional stability analysis, mainly focuses on the study of the system’s trajectory to converge into an equilibrium point or some other nominal trajectory. Compared to these studies, incremental stability proposes a stronger notion of stability, denoting convergence of arbitrary trajectories towards each other rather than to a particular point [4]. As a result, it has gained significant attention for its broad applicability, including nonlinear analog circuit modeling [8], cyclic feedback system synchronization [14], symbolic model development [28, 42, 17, 16, 30], stability of interconnected systems [20], oscillator synchronization [32], and complex network analysis [29].
@@
Several promising tools have been developed in the last decades to analyze the property of incremental stability. Contraction analysis [22] and convergent dynamics [27] have emerged as promising tools to analyze incremental stability properties for nonlinear systems. However, traditional stability analysis is done using comparison functions, which leads to the notion of invariance of the system with respect to change in coordinates [13]. This leads the research community to find the most powerful tool to analyze incremental stability, namely incremental Lyapunov functions [4, 35, 39]. Later, these tools have also been extended to analyze the incremental stability of a wide class of systems, such as nonlinear systems [4], stochastic systems [7], hybrid dynamical systems [7], time-delayed systems [9], and interconnected switched systems [10].
@@
The incremental stability property has been identified as a promising tool for the construction of finite abstraction of nonlinear control systems [12]. Hence, it is important to design or synthesize controllers that enforce incremental stability. However, designing such controllers has been challenging over the years as most of the existing approaches to ensure incremental stability [41, 43, 15] rely on the complete knowledge of the dynamics of the systems. In particular, these systems should possess specific structures (for example, control-affine systems). However, in practice, uncertainties often lead to incomplete or inaccurate system models, which ultimately lead to the fact that these controllers will not be effective in real-world scenarios. In addition, in some cases, the system models are so complex that even model derivation is a tedious job. In these scenarios, to address the limitations posed by unknown dynamics or modeling errors, learning-based approaches have gained significant attention. The authors in [33] employ the Gaussian process to collect data from the system and develop a backstepping-like controller, assuming the system follows a control-affine structure. Similarly, the latest approach for designing a controller to achieve incremental ISS [38] utilizes a data-driven technique, where data is extracted from system trajectories. However, this method relies on the fact that the system possesses a polynomial-type structure.
@@
Among the learning-based approaches for verification and controller synthesis, deep learning-based techniques have gained prominence to estimate system dynamics or synthesize controllers for different task specifications alongside constructing Lyapunov or barrier functions to guarantee stability or safety. The existing literature [1, 3, 34, 25] addresses the issue of safety and stability while eliminating the need for explicit knowledge of the system dynamics. Leveraging the universal approximation capabilities of neural networks, functions can be directly synthesized. The key challenge to use neural networks as Lyapunov functions or controllers is to provide a formal guarantee, as the training relies on discrete samples, which cover only a limited part of the continuous state space.
@@
Contributions: The previous work of the author proposes a formally verified neural network controller to ensure incremental input-to-state stability for unknown discrete-time systems [6]. This represents a significant advancement beyond the formal verification of neural Lyapunov functions for unknown systems [5], as it not only ensures stability but also provides a systematic approach to controller design with rigorous verification guarantees. In this work, we aim to synthesize the controller for unknown continuous-time systems such that the closed-loop system exhibits the behavior of incremental input-to-state stability. We propose the notion of an incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF) for continuous-time systems and prove that the existence of such a function under the controller leads the closed-loop system to be incrementally stable. In this work, as we deal with general nonlinear systems, we are using data-driven techniques to synthesize the controller, therefore, we need the compactness assumption on the state-space. As the controller and $\delta$-ISS-CLF will be formulated over finite data, we need to ensure forward invariance of the state-space i.e., starting from any initial condition within the state-space, the trajectory of the system always lies inside the state-space. To ensure forward invariance, we utilize the notion of control barrier function (CBF) to make the state-space forward-invariant. We simultaneously leverage the conditions of $\delta$-ISS-CLF with the CBF condition imposed on the controller as a robust optimization problem. The system dynamics being unknown, we propose the controller and the $\delta$-ISS-CLF as neural networks and formulate a scenario convex problem corresponding to the robust optimization problem using some Lipschitz continuity assumptions. The conditions of Lyapunov and barrier are used to formulate the loss function for training neural networks. Finally, we propose a validity condition that gives a formal guarantee that the trained neural controller can make the system incrementally stable. We validate the effectiveness of our approach by applying it to multiple case studies - the first one is a scalar system with a non-affine structure, the second example is a one-link manipulator system, the third system is a nonlinear Moore-Grietzer model of the jet engine and the final one is a rotating rigid spacecraft model.
The key contributions of this paper are highlighted below.
  • 1.
    For the first time, we introduce the notion of incremental input-to-state stable control Lyapunov functions for continuous-time systems.
  • 2.
    Given that the proposed approach requires data for training neural networks, it necessitates reliance on compact sets. To address this, we establish conditions for incremental input-to-state stable control Lyapunov functions under compact sets, ensuring incremental input-to-state stability for systems evolving in compact sets.
  • 3.
    A novel training framework is introduced to concurrently synthesize a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parameterized as neural networks, for unknown continuous-time systems.
  • 4.
    Previous studies to design the controller to enforce incremental stability required that the system should possess either a parametric strict-feedback type structure [31] or a strict-feedback type structure [43] or Hamiltonian systems [15]. In this work, we do not require any knowledge of the system dynamics or underlying structure. Thus, we propose to synthesize a controller for any general system without any assumptions.
  • 5.
    As the controller is a neural network, we can control the system input of higher dimensions with a smaller number of external inputs to the controller. Section 5.5 details the implications of reduced control input space for an incrementally stable system.
  • 6.
    For the continuous-time system, the derivative of the neural $\delta$-ISS-CLF should have Lipschitz continuity. In this work, we propose that the derivative of a multi-layer feedforward neural network can be bounded by another neural network, provided the activation functions used are Lipschitz continuous. Using the standard Lipschitz constant estimation procedure of a neural network, we ensure a bound on the Lipschitz constant of the derivative of the neural $\delta$-ISS-CLF.
@@
The organization of the paper is as follows. In Section 2, we recall the notion of incremental input-to-state stability and introduce the notion of incremental input-to-state stable control Lyapunov function. In Section 3, we propose a learning framework that jointly synthesizes a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parametrized as neural networks with the system dynamics being unknown. We pose the problem of synthesizing the controller as an optimization problem and solve it with the help of neural network training, as discussed in Section 4. Finally, in Section 5, we validate the proposed approach by showcasing multiple case studies and providing a general discussion on the usefulness of the results.
@@ 2.1 Notations
The symbols ${\mathbb{N}}$, ${\mathbb{N}}_{0}$, ${\mathbb{R}}$, ${\mathbb{R}}^{+}$, and ${\mathbb{R}}_{0}^{+}$ denote the set of natural, nonnegative integer, real, positive real, and nonnegative real numbers, respectively. A vector space of real matrices with $n$ rows and $m$ columns is denoted by ${\mathbb{R}}^{n\times m}$. A column vector with $n$ rows is represented by ${\mathbb{R}}^{n}$. The Euclidean norm is represented using ||$\cdot$|| ⋅ |. Given a function $\varphi:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}^{m}$, its sup-norm (possibly $\infty$-norm) is given by $\lVert\varphi\rVert=\sup\{|\varphi(k)|:k\in{\mathbb{R}}_{0}^{+}\}$. For $a,b\in{\mathbb{R}}_{0}^{+}$ with $a\leq b$, the closed interval in ${\mathbb{N}}_{0}$ is denoted as $[a;b]$. A vector $x\in\mathbb{R}^{n}$ with entries $x_{1},\ldots,x_{n}$ is represented as $[x_{1},\ldots,x_{n}]^{\top}$, where $x_{i}\in\mathbb{R}$ denotes the $i$-th element of the vector and $i\in[1;n]$. A diagonal matrix in ${\mathbb{R}}^{n\times n}$ with nonnegative entries is denoted by $\mathcal{D}_{\geq 0}^{n}$. Given a matrix $M\in{\mathbb{R}}^{n\times m}$, $M^{\top}$ represents the transpose of matrix $M$. A continuous function $\alpha:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be class $\mathcal{K}$, if $\alpha(s)>0$ for all $s>0$, strictly increasing and $\alpha(0)=0$. It is class $\mathcal{K}_{\infty}$ if it is class $\mathcal{K}$ and $\alpha(s)\rightarrow\infty$ as $s\rightarrow\infty$. A continuous function $\beta:{\mathbb{R}}_{0}^{+}\times{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a class $\mathcal{KL}$ if $\beta(s,t)$ is a class $\mathcal{K}$ function with respect to $s$ for all $t$ and for fixed $s>0$, $\beta(s,t)\rightarrow 0$ if $t\rightarrow\infty$. For any compact set $\mathcal{C},\partial\mathcal{C}$ and $int(\mathcal{C})$ denote the boundary and interior of the set $\mathcal{C}$, respectively. The gradient of a function $f:{\mathbf{X}}\rightarrow{\mathbb{R}}$ is denoted by $\nabla f(x)$. A function $f:{\mathbf{X}}\rightarrow{\mathbb{R}}$ is said to $\vartheta$-smooth if for all $x,y\in{\mathbf{X}},|\nabla f(x)-\nabla f(y)|\leq\vartheta|x-y|$. The partial differentiation of a function $f:{\mathbf{X}}\times\hat{{\mathbf{X}}}\rightarrow{\mathbb{R}}$ with respect to a variable $x\in{\mathbf{X}}$ is given by $\frac{\partial f}{\partial x}$.
@@ 2.2 Incremental Stability for Continuous-time systems
We consider a nonlinear continuous-time control system (ct-CS) represented by the tuple $\Xi=({\mathbf{X}},{\mathbf{U}},f)$, where the state of the system is $\mathsf{x}(t)\in{\mathbf{X}}\subseteq{\mathbb{R}}^{n}$ and input to the system is $\mathsf{u}(t)\in{\mathbf{U}}\subset{\mathbb{R}}^{m}$ at time $t\geq 0$. The state dynamics is described using the following differential equation
$\dot{\mathsf{x}}=f(\mathsf{x},\mathsf{u}).$ (1)
Next, the notion of the closed-loop continuous-time control system under feedback controller $g$ is defined, which is represented as $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$, where ${\mathbf{X}}\subseteq{\mathbb{R}}^{n}$ is the state-space of the system, ${\mathbf{U}}\subseteq{\mathbb{R}}^{m}$ is the internal input set of the system, ${\mathbf{W}}\subseteq{\mathbb{R}}^{p}$ is the external input set of the system, $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ and $f:{\mathbf{X}}\times{\mathbf{U}}\rightarrow{\mathbf{X}}$ are maps describing the state evolution as:
@@
Definition 2.1 ($\delta$-ISS [4])
@@
$|\mathsf{x}_{x,\mathsf{w}}(t)-\mathsf{x}_{\hat{x},\hat{\mathsf{w}}}(t)|\leq\beta(|x-\hat{x}|,t)+\gamma(\lVert\mathsf{w}-\hat{\mathsf{w}}\rVert).$ (3)
If $\mathsf{w}=\hat{\mathsf{w}}$, one can recover the notion of incremental global asymptotic stability as defined in [4].
Next, we introduce the concept of incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF) for continuous-time systems, which defines the sufficient conditions to ensure incremental stability of the closed-loop system with the controller. The sufficient conditions are defined next:
@@
A smooth function $V:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a $\delta$-ISS control Lyapunov function for closed-loop system $\Xi_{g}=({\mathbb{R}}^{n},{\mathbb{R}}^{p},{\mathbb{R}}^{m},f,g)$, if there exist a controller $g:{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\rightarrow{\mathbb{R}}^{m}$, class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\sigma$ and a constant $\kappa\in{\mathbb{R}}^{+}$ such that:
@@
    for all $x,\hat{x}\in{\mathbb{R}}^{n}$ and for all $w,\hat{w}\in{\mathbb{R}}^{p}$,
    $\frac{\partial V}{\partial x}f(x,g(x,w))+\frac{\partial V}{\partial\hat{x}}f(\hat{x},g(\hat{x},\hat{w}))\leq-\kappa V(x,\hat{x})+\sigma(|w-\hat{w}|)$.
The following theorem describes $\delta$-ISS of the closed-loop system in terms of the existence of $\delta$-ISS control Lyapunov function.
@@
The closed-loop continuous-time control system (2) will be incrementally input-to-state stable with respect to external input ${\mathsf{w}}$ if there exists a $\delta$-ISS control Lyapunov function satisfying the conditions of Definition 2.2.
@@
The proof can be done using a similar approach as in [4, Theorem 2]. $\square$
In this work, we aim to solve the problem of incremental stability using a data-driven approach. This requires working with compact sets; hence, we introduce the notion of $\delta$-ISS-CLF for compact sets. To do so, we recall the notion of control forward invariance.
Definition 2.4 (Control Forward Invariant Set [36])
A set ${\mathbf{X}}$ is said to be control forward invariant with respect to the system (1), if for every initial condition $x\in{\mathbf{X}}$, there exists a control input sequence $\mathsf{u}$ such that $\mathsf{x}_{x,\mathsf{u}}(t)\in{\mathbf{X}},\forall t\geq 0$. Now, if there exists a controller $g:{\mathbf{X}}\rightarrow{\mathbf{U}}$ such that the set ${\mathbf{X}}$ becomes control forward invariant, then the controller is said to be a forward-invariant controller.
@@
A smooth function $V:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a $\delta$-ISS control Lyapunov function for closed-loop system $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$ in (2), where ${\mathbf{X}},{\mathbf{W}}$ are compact sets, if there exist a forward invariant controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$, class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\sigma$ and a constant $\kappa\in{\mathbb{R}}^{+}$ such that:
@@
    for all $x,\hat{x}\in{\mathbf{X}}$ and for all $w,\hat{w}\in{\mathbf{W}}$,
    $\frac{\partial V}{\partial x}f(x,g(x,w))+\frac{\partial V}{\partial\hat{x}}f(\hat{x},g(\hat{x},\hat{w}))\leq-\kappa V(x,\hat{x})+\sigma(|w-\hat{w}|)$.
@@
The closed-loop ct-CS (2) is said to be incrementally input-to-state stable within the state space ${\mathbf{X}}$ with respect to the external input ${\mathsf{w}}$, if there exists a $\delta$-ISS control Lyapunov function under the forward invariant controller $g$ as defined in Definition 2.5.

### v2

# Neural Controller for Incremental Stability of Unknown Continuous-time Systems

@@ Abstract
This work primarily focuses on synthesizing a controller that guarantees an unknown continuous-time system to be incrementally input-to-state stable ($\delta$-ISS). In this context, the notion of $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for the continuous-time system is introduced. Combined with the controller, the $\delta$-ISS-CLF guarantees that the system is incrementally stable. As the paper deals with unknown dynamical systems, the controller as well as the $\delta$-ISS-CLF are parametrized using neural networks. The data set used to train the neural networks is generated from the state space of the system by proper sampling. Now, to give a formal guarantee that the controller makes the system incrementally stable, we develop a validity condition by having some Lipschitz continuity assumptions and incorporate the condition into the training framework to ensure a provable correctness guarantee at the end of the training process. Finally, we demonstrate the effectiveness of the proposed approach through several case studies: a scalar system with a non-affine, non-polynomial structure, a one-link manipulator system, a nonlinear Moore-Greitzer model of a jet engine, a magnetic levitator system and a rotating rigid spacecraft model.
@@ 1 Introduction
Input-to-state stability has been a promising tool for analyzing and studying robust stability questions for nonlinear control systems. This traditional stability analysis mainly focuses on the study of the system’s trajectory to converge into an equilibrium point or some other nominal trajectory. Compared to these studies, incremental stability proposes a stronger notion of stability, denoting convergence of arbitrary trajectories towards each other rather than to a particular point [angeli2002lyapunov]. As a result, it has gained significant attention for its broad applicability, including nonlinear analog circuit modeling [analog], cyclic feedback system synchronization [cyclic_feedback], symbolic model development [symbolic1, zamani2017towards, jagtap2020symbolic, jagtap2017quest, sadek2023compositional], stability of interconnected systems [DD-Stability], oscillator synchronization [inter_osci], and complex network analysis [synchComplex].
@@
Several promising tools have been developed in the last decades to analyze the property of incremental stability. Contraction analysis [Contraction, sun2021learning, tsukamoto] and convergent dynamics [Conv_dyn] have emerged as promising tools to analyze incremental stability properties for nonlinear systems. However, traditional stability analysis is done using comparison functions, which leads to the notion of invariance of the system with respect to change in coordinates [grune1999asymptotic]. This leads the research community to find the most powerful tool to analyze incremental stability, namely incremental Lyapunov functions [angeli2002lyapunov, DT-ISS_prop, zamani2011lyapunov]. Later, these tools have also been extended to analyze the incremental stability of a wide class of systems, such as nonlinear systems [angeli2002lyapunov], stochastic systems [biemond2018incremental], hybrid dynamical systems [biemond2018incremental], time-delayed systems [chaillet2013razumikhin], and interconnected switched systems [dey2023incremental].
@@
Incremental stability is a powerful tool for the construction of a finite abstraction of nonlinear control systems [girard2009approximately], which motivates the design of controllers that enforce it. However, most of the existing approaches to design controllers ensuring incremental stability [zamani2011backstepping, zamani2013backstepping, jagtap2017backstepping] rely on the complete knowledge of the dynamics of the systems, making them unsuitable when uncertainties or modelling errors are present in real-world scenarios. To address these limitations posed by unknown dynamics or modelling errors, data-driven and learning-based approaches have gained significant attention. The authors in [sangeerth2025controller, sundarsingh2024backstepping] employ the Gaussian process to collect data from the system and develop a backstepping-like controller. However, this method assumes the system to be control-affine and then designs the controller. Compared to it, a very recent approach for designing a controller to achieve incremental ISS [zaker2024controller] utilizes a data-driven technique, but is only limited to polynomial-type structured dynamics, where data is extracted from system trajectories. However, both these works need the system to be identified first before the control design step, which is indeed challenging.
@@
Among the learning-based approaches for verification and controller synthesis, deep learning-based techniques have gained prominence to estimate system dynamics or synthesize controllers for different task specifications alongside constructing Lyapunov or barrier functions to guarantee stability or safety. The existing literature [formal_nn_Lyapunov, Neural_CBF, tayal2024learning] addresses the issue of safety and stability while eliminating the need for explicit knowledge of the system dynamics. Leveraging the universal approximation capabilities of neural networks, functions can be directly synthesized. Some recent works have used neural networks adaptively to learn the controller in an iterative fashion [shen2024neural] while data-driven control via Q-learning [shen2024event] has been employed to solve tracking problems as well. However, the key challenge to use neural networks as Lyapunov functions or controllers is to provide a formal guarantee, as the training relies on discrete samples, which cover only a limited part of the continuous state space. While falsification-based approaches [chang2019neural] have been proposed to mitigate this by adaptively sampling “difficult” regions, they provide only probabilistic confidence; in contrast, we try to enforce deterministic satisfaction of $\delta$-ISS conditions through constrained optimization. The CEGIS-based approach [abate2020formal] has been a powerful tool as well to verify neural Lyapunov function, but it works in a post-hoc verification scenario, while in this work, we incorporate verification within the training procedure without the need for any other solver.
@@
Contributions: Building on our earlier work on formally verified neural network controllers for unknown discrete-time systems [basu2025controller], we extend the framework to unknown continuous-time systems. We propose the notion of an incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF) for continuous-time systems and prove that the existence of such a function under the controller leads the closed-loop system to be incrementally stable. For a general unknown system, we propose data-driven techniques to synthesize $\delta$-ISS-CLF and the controller jointly and hence the notion of $\delta$-ISS-CLF over compact sets is proposed while ensuring forward invariance by leveraging the notion of CBF as seen in [basu2025controller]. We unify the $\delta$-ISS-CLF and CBF conditions in a robust optimization program (ROP) and propose the controller and the $\delta$-ISS-CLF as neural networks, as system dynamics is unknown and only a black-box model is provided, and formulate a scenario convex problem corresponding to the ROP using some Lipschitz continuity assumptions. We propose a validity condition that gives a formal guarantee that the neural controller can make the system incrementally stable and incorporate it in the training framework to avoid any posteriori verification. We validate the effectiveness of our approach by applying it to multiple case studies.
The major contribution of this article over the discrete-time version is twofold. First, the paper highlights the notion of $\delta$-ISS-CLF for continuous time systems over compact sets, which is completely new compared to the existing literature. Second, the derivative of $\delta$-ISS-CLF to ensure $\delta$-ISS needs to be bounded to provide a deterministic guarantee over the controller, for which it requires the derivative of a neural network to be bounded by some Lipschitz constant. In this work, we propose a novel loss function that ensures the derivative of a multi-layer feedforward neural network is also Lipschitz continuous.
@@
The organization of the paper is as follows. In Section 2, we recall the notion of incremental input-to-state stability and introduce the notion of incremental input-to-state stable control Lyapunov function for compact sets. In Section 3, we propose a learning framework that jointly synthesizes a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parametrized as neural networks, posing the problem of synthesizing the controller as an optimization problem and solve it with the help of neural network training, as discussed in Section 4. Finally, in Section 5, we validate the proposed approach by showcasing multiple case studies and providing a general discussion on the usefulness of the results.
@@ 2.1 Notations
The symbols $\mathbb{N}$}, ${\mathbb{N}}_{0}$, $\mathbb{R}$}, $\mathbb{R}$$^{+}$, and ${\mathbb{R}}_{0}^{+}$ denote the set of natural, nonnegative integers, real, positive real, and nonnegative real numbers, respectively. A vector space of real matrices with $n$ rows and $m$ columns is denoted by ${\mathbb{R}}^{n\times m}$. The column vector space with $n$ rows is represented by ${\mathbb{R}}^{n}$. The Euclidean norm is represented using $|\cdot|$. Given a function $\varphi:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}^{m}$, its sup-norm (possibly $\infty$-norm) is given by $\lVert\varphi\rVert=\sup\{|\varphi(k)|:k\in{\mathbb{R}}_{0}^{+}\}$. For $a,b\in{\mathbb{N}}_{0}$ with $a\leq b$, the closed interval in ${\mathbb{N}}_{0}$ is denoted as $[a;b]$. A vector $x\in\mathbb{R}^{n}$ with entries $x_{1},\ldots,x_{n}$ is represented as $x_{1},\ldots,x_{n}]^{\top}$, where $x_{i}\in\mathbb{R}$ denotes the $i$-th element of the vector and $i\in[1;n]$. The space of diagonal matrices in ${\mathbb{R}}^{n\times n}$ with nonnegative entries is denoted by $\mathcal{D}_{\geq 0}^{n}$. Given a matrix $\in{\mathbb{R}}^{n\times m}, MM^{\top}$ represents the transpose of matrix $M$. A continuous function $\alpha:{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be class $\mathcal{K}$, if $\alpha(s)>0$ for all $s>0$, strictly increasing and $\alpha(0)=0$. It is class $\mathcal{K}_{\infty}$ if it is class $\mathcal{K}$ and $\alpha(s)\rightarrow\infty$ as $s\rightarrow\infty$. A continuous function $\beta:{\mathbb{R}}_{0}^{+}\times{\mathbb{R}}_{0}^{+}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a class $\mathcal{KL}$ if $\beta(s,t)$ is a class $\mathcal{K}$ function with respect to $s$ for all $t$ and for fixed $s>0$, $\beta(s,t)\rightarrow 0$ if $t\rightarrow\infty$. For any compact set $\mathcal{C},\partial\mathcal{C}$ and $int(\mathcal{C})$ denote the boundary and interior of the set $\mathcal{C}$, respectively. The gradient of a function $f:{\mathbf{X}}\rightarrow{\mathbb{R}}$ is denoted by $\nabla f(x)$. A function $f:{\mathbf{X}}\rightarrow{\mathbb{R}}$ is said to $\vartheta$-smooth if for all $x,y\in{\mathbf{X}},|\nabla f(x)-\nabla f(y)|\leq\vartheta|x-y|$. The partial differentiation of a function $f:{\mathbf{X}}\times\hat{{\mathbf{X}}}\rightarrow{\mathbb{R}}$ with respect to a variable $x\in{\mathbf{X}}$ is given by $\frac{\partial f}{\partial x}$.
@@ 2.2 Incremental Stability for Continuous-time systems
We consider a non-linear continuous-time control system (ct-CS) represented by the tuple $\Xi=({\mathbf{X}},{\mathbf{U}},f)$ and describe it using the differential equation as
$\dot{\mathsf{x}}=f(\mathsf{x},\mathsf{u}),$ (1)
where $\mathsf{x}(t)\in{\mathbf{X}}\subseteq{\mathbb{R}}^{n}$ is the state of the system and $\mathsf{u}(t)\in{\mathbf{U}}\subseteq{\mathbb{R}}^{m}$ is the input to the system at time $t\geq 0$, the function $f:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{n}$ is assumed to be locally Lipschitz continuous to ensure the existence and uniqueness of the solution [khalil2002nonlinear].
Next, the notion of the closed-loop continuous-time control system under a Lipschitz continuous feedback controller $g$ is defined, which is represented as $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$, where ${\mathbf{X}}\subseteq{\mathbb{R}}^{n}$ is the state-space of the system, ${\mathbf{U}}\subseteq{\mathbb{R}}^{m}$ is the internal input set of the system, ${\mathbf{W}}\subseteq{\mathbb{R}}^{p}$ is the external input set of the system, $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ and $f:{\mathbf{X}}\times{\mathbf{U}}\rightarrow{\mathbb{R}}^{n}$ are maps describing the state evolution as:
@@
Definition 2.1 ($\delta$-ISS [angeli2002lyapunov])
@@
[…]
If $\mathsf{w}=\hat{\mathsf{w}}=0$, one can recover the notion of incremental global asymptotic stability as defined in [angeli2002lyapunov]. Next, we introduce the concept of incremental input-to-state stable control Lyapunov function ($\delta$-ISS-CLF) for continuous-time systems.
@@
A smooth function $V:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a $\delta$-ISS control Lyapunov function ($\delta$-ISS-CLF) for closed-loop system $\Xi_{g}=({\mathbb{R}}^{n},{\mathbb{R}}^{p},{\mathbb{R}}^{m},f,g)$ with the feedback controller $g:{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\rightarrow{\mathbb{R}}^{m}$, if there exist class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\sigma$ and a constant $\kappa\in{\mathbb{R}}^{+}$ such that:
@@
[…]
Intuitively, the $\delta$-ISS-CLF quantifies the contraction behavior between any two system trajectories. The first condition enforces positive semidefiniteness with respect to the state difference, while the second condition guarantees exponential decay of this difference along the closed-loop trajectories. It should be carefully noted that the proposed $\delta$-ISS-CLF notion differs from classical incremental Lyapunov functions by explicitly incorporating a feedback control law. This allows the function to serve as a constructive tool for controller synthesis. Now, the following theorem describes $\delta$-ISS of the closed-loop system in terms of the existence of a $\delta$-ISS-CLF.
@@
The closed-loop continuous-time control system (2) will be $\delta$-ISS with respect to external input $\mathsf{w}$} if there exists a $\delta$-ISS-CLF satisfying the conditions of Definition $2.$
@@
The proof follows similar to that of Theorem 2 of [angeli2002lyapunov]. $\square$
In this work, we aim to solve the problem of incremental stability using a data-driven approach. This requires the collection of data from compact sets; hence, we need the notion of $\delta$-ISS-CLF for compact sets. To do so, we first recall the notion of control forward invariance.
Definition 2.4 (Control Forward Invariant Set [liu2019compositional])
A set $\mathbf{X}$} is said to be control forward invariant with respect to the system (2) if for any $(x,w)\in{\mathbf{X}}\times{\mathbf{W}}$, there exists some control input $u\in{\mathbf{U}}$ such that $f(x,u)\in{\mathbf{X}}$. Now, if there exists a controller $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ such that the set $\mathbf{X}$} becomes control forward invariant, then the controller is said to be a forward invariant controller corresponding to the forward invariant set $\mathbf{X}$} with respect to the external input set $\mathbf{W}$}.
@@
A smooth function $V:{\mathbf{X}}\times{\mathbf{X}}\rightarrow{\mathbb{R}}_{0}^{+}$ is said to be a $\delta$-ISS control Lyapunov function for closed-loop system $\Xi_{g}=({\mathbf{X}},{\mathbf{W}},{\mathbf{U}},f,g)$ in (2), where ${\mathbf{X}},{\mathbf{W}}$ are compact sets and $g:{\mathbf{X}}\times{\mathbf{W}}\rightarrow{\mathbf{U}}$ be a forward invariant feedback controller, if there exist class $\mathcal{K}_{\infty}$ functions $\alpha_{1},\alpha_{2},\sigma$ and a constant $\kappa\in{\mathbb{R}}^{+}$ such that:
@@
[…]
@@
The closed-loop ct-CS (2) is said to be $\delta$-ISS within the compact state space $\mathbf{X}$} with respect to the external input $\mathsf{w}$}, if there exists a $\delta$-ISS-CLF under the forward invariant controller $g$ as defined in Definition 2.5.